OpenStudy (anonymous):

Set up the Surface area Double integration f(x,y)=cos(x^2+y^2) R={(x,y), x^2+y^2<=pi/2}

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi/2}rcos(r^2)drd \theta\]

OpenStudy (anonymous):

correct?

OpenStudy (lukecrayonz):

yes

OpenStudy (anonymous):

actually it'd be pi/4 right? the diameter is pi/4

OpenStudy (anonymous):

pi/2

OpenStudy (anonymous):